Use the method of contradiction to prove that √5 is not a rational number.
1
Expert's answer
2020-05-25T21:11:29-0400

Solution.

Assume that √5 is a rational number. Suppose there exists a rational number whose square is 5. This number can be given as an unbreakable fraction "\dfrac{m}{n}" where "m,n -" are natural numbers.

Then "\sqrt{5}=\dfrac{m}{m}, 5=\dfrac{m^2}{n^2}, m^2=5n^2."

"m^2" is divisible by 5, so "m" is divisible by 5. Then "m^2" is divisible by 25. If so, then "n^2" is divisible by 5, and therefore "n" is divisible by 5. It turns out that the fraction "\dfrac{m}{n}" is short. We came to a contradiction of the original judgment (that the number is rational). Therefore, the number "\sqrt{5}" is irrational.

Answer: the number  "\sqrt{5}" is irrational.